25cm^3 of a solution of barium chloride reacted with 20cm^3 of a solution of 0.05mol/dm^3 sulphuric acid to form barium sulphate and hydrochloric acid.
A) write a balanced chemical equation for the reaction
B) calculate the concentration of barium chloride solution
C) name the precipitate formed in the reaction
1
Expert's answer
2020-10-14T05:45:44-0400

A) BaCl2+H2SO4=BaSO4+2HCl.

B) C1V1=C2V2,

C1=C2V2/V1,

C(BaCl2)=0.05×20/25=0.04(mol/dm3).

C) Barium sulphate.